OpenStudy (yanasidlinskiy):

Write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. Vertex at (-5,0) and co-vertex at (0,4).

OpenStudy (yanasidlinskiy):

@mathmale Can you help me??

OpenStudy (yanasidlinskiy):

@jim_thompson5910

OpenStudy (mathmale):

I'd strongly suggest that you graph this ellipse as best you can. Remember that this graph will have symmetry: vertical and horizontal. Show the vertex and the co-vertex.

OpenStudy (mathmale):

would you say that this ellipse is horizontal or vertical? How would you know?

OpenStudy (yanasidlinskiy):

I don't really know cuz I haven't learned these ellipses at all.. But let me post the choices..hang on..

OpenStudy (yanasidlinskiy):

a.\[\frac{ x ^{2} }{ 25 }+\frac{ y^2 }{ 25 }\] =1 b.\[\frac{ x^2 }{ 16 }+\frac{ y^2 }{ 25 }\] =1 c. \[\frac{ x^2 }{ 5 } +\frac{ y^2 }{ 4 }=1\] d. \[\frac{ x^2 }{ 4 }+\frac{ y^2 }{ 5 }=1\]

OpenStudy (mathmale):

Yana, what kind of learning materials are you working with? Have you a textbook, online materials, or do you actually attend classroom lectures? I hate to see you struggling in the dark. You need, as I did, some reference so that you could learn the basics of conic sections (including ellipses). I will sketch this ellipse based upon the 2 points given:

OpenStudy (mathmale):



OpenStudy (mathmale):

Yana, the info given seems to indicate that the center of this ellipse is (0,0) (the origin). If that is so, then the equation of this ellipse has the form\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] Have you seen this before? If so, where?

OpenStudy (mathmale):

If you'll look at my diagram again, you'll see the values of a and b. Please substitute these into the equation I've just typed out. When you're done, you'll have the equation you wanted for this particular ellipse.

OpenStudy (yanasidlinskiy):

I have not seen that before....hang on let me get the equation and tall me if im right..

OpenStudy (yanasidlinskiy):

Just give me some time....real quick:)

OpenStudy (yanasidlinskiy):

ok...is it c??

OpenStudy (yanasidlinskiy):

\[\frac{ x^2 }{ 5^2 }+\frac{ y^2 }{ 4^2 }=1\]??

OpenStudy (yanasidlinskiy):

wait i kind of messed up..